### v1

@@ Abstract
We show that a Lie group $G$ admitting a bi-invariant distance must be the product $G=H\times K$ of an abelian group $H$ and a compact group $K$ with discrete center. Moreover, the distance in $G$ must come from the infima of lengths of paths for a unique infinitesimal metric (a Finsler norm) defined in the Lie algebra of $G$. From this we derive the distance minimizing paths which are left or right translations of one-parameter groups (though these are not the unique minizing paths if the norm is not smooth or strictly convex). Then we introduce a notion of sectional curvature $\sec(\pi)$ for a bi-invariant distance, following Milnor’s ideas, and we show that this curvature is bounded and non-negative, and it is null when the $2$-plane $\pi$ is an abelian Lie subalgebra of $Lie(G)$. We show that when the distance is strictly convex, our sectional curvature vanishes if and only if the $2$-plane is abelian. We give finer characterizations for the case of vanishing curvature, for the case of non-strictly convex norms. [^4] Keywords and phrases: bi-invariant metric, curvature, Finsler metric, geodesic, Lie group, one-parameter group, skew-Hermitian matrix, unitary group, unitarily invariant norm
@@ 1 Introduction
This paper is at large, a study of the geometry of a Lie group with a bi-invariant distance $(G,\operatorname{dist})$. To begin with, we consider a Finsler norm in the Lie algebra $\operatorname{\mathfrak{g}}$ of $G$. More precisely: a subadditive and positively homogeneous map $|\cdot|:\operatorname{\mathfrak{g}}\to\mathbb{R}_{\geq 0}$ is a Finsler norm if its non-degenerate (i.e. $|x|=0$ implies $x=0$). We do not require smoothness neither full homogeneity. By requiring $x_{n}\to x$ if $|x_{n}-x|\to 0$, the Finsler norm induces the vector space topology in $\operatorname{\mathfrak{g}}$. We say that the Finsler norm is $\operatorname{Ad}$-invariant if $|\operatorname{Ad}_{g}x|=|x|$ for any $g\in G$ and any $x\in\operatorname{\mathfrak{g}}$. By left or right translation, this gives a way to measure tangent vectors in $TG$. Therefore the $\operatorname{Ad}$-invariant Finsler norm induces in the group $G$ a bi-invariant distance (which might not be reversible: $d(x,y)=d(x,y)$ if and only if the Finsler norm is fully homogeneous) by taking the infima of the legths of paths joining $x,y$. One of the pleasant features of dealing with such generality is that we are dealing with convex bodies around the origin of $\operatorname{\mathfrak{g}}$, which are balanced if and only if the norm is fully homogeneous. Another important point to stress is that any bi-invariant distance on a locally Euclidean topological group $G$ can be obtained like this, see Theorem (A) below.
This paper is a natural continuation of [26], where certain key properties of $\operatorname{Ad}$-invariant Finsler norms in the space of skew-Hermitian matrices were established. The purpose of this paper is to obtain a reasonable notion of curvature for groups with bi-invariant distances, following the guidelines of metric geometry where curvature is a measure of the distortion of the distance, with respect to a certain model space. More to the point, following Milnor [29, p. 101] we define a suitable notion of the curvature of $G$ in the direction of vectors $x,y$, as done for instance in [3] in a different setting. It builds on our previous work on skew-Hermitian matrices [26], and also gives some insight to the elliptic geometry of the unitary groups with bi-invariant metrics, in opposition to the hyperbolic geometry studied in [3]. In terms of complexifications, if $G$ is a compact complex Lie group with Lie algebra $\operatorname{\mathfrak{g}}$ and $\operatorname{\mathfrak{k}}$ is its real compact form, where $\operatorname{\mathfrak{g}}=\operatorname{\mathfrak{k}}\oplus i\operatorname{\mathfrak{k}}$, then ours is the study of the geometry of $G_{\mathbb{R}}$ with a (non-necessarily Riemannian) bi-invariant distance, in terms of the properties of $\operatorname{\mathfrak{k}}$. On the other hand, the results in [3] are related to the geometry of the homogeneous space $G/G_{\mathbb{R}}\simeq i\mathbb{\operatorname{\mathfrak{k}}}$.
Since a Lie group $G$ with a bi-invariant Riemannian metric can be fully characterized [30] as the product of a vector space and a compact Lie group $K$, the natural question that arose during this research was whether it was possible to give such characterization knowing only the existence of a bi-invariant (non-necessarily Riemannian) distance in $G$. We could solve this problem in the affirmative: the main result of Section 2 is the following
@@
If $(G,\operatorname{dist})$ is a connected, locally compact, locally contractible topological group with a bi-invariant intrinsic distance, then $G=H\times K$ is the product of an abelian connected Lie group $H$ and a connected compact Lie group with discrete center $K$.
Locally, the distance between $e^{sx}$ and $e^{sy}$ can be computed as the norm of the Baker-Campbell-Hausdorff series of $e^{sy}e^{-sx}$, that is $\operatorname{dist}(e^{sx},e^{sy})=BCH(sy,-sx)$. In particular if $[x,y]=0$ then $\operatorname{dist}(e^{sx},e^{sy})=|sy-sx|$ for small values of $s$. This was stablished in previous research, see for instance [24].
In Secion 3.3 we take the approach of J. Milnor [29, p.101], where sectional curvature of the $2$-plane $\pi\subset\mathfrak{g}$ can be defined/recovered from the metric as follows:
Using the formulas for the distance involving the BCH formula, we then show that the sectional curvature of $G$ can be expressed in purely metric terms, and we prove the following:
@@
For $x,y\in\operatorname{\mathfrak{g}}$ and any $\operatorname{Ad}$-invariant norm in $\operatorname{\mathfrak{g}}$ we have that
$S(x,y)=6|y-x|\lim\limits_{r\to 0^{+}}\frac{\operatorname{dist}(e^{r^{2}x},e^{r^{2}y})-r\operatorname{dist}(e^{rx},e^{ry})}{r^{4}}=\frac{-|y-x|}{4}\max\limits_{\varphi\in N_{y-x}}\varphi([x,[x,y-x]])\geq 0.$
where $N_{y-x}$ denotes the set of all unit norm linear functionals such that $\varphi(y-x)=|y-x|$.
The remarkable observation of Milnor was that, if the distance comes from a Riemannian metric, and $x,y$ are normalized, then $S(x,x+y)=\langle R(x,y)y,x\rangle$ where $R$ is the curvature tensor of the Levi-Civita connection. What we give is a characterization in terms of norming functionals (the subdifferential of the norm).
Finally, in the last section of the paper we study the flatness condition in terms of norming functionals and the form $S$, and in Theorem 3.28, equipped with the machinery of a root system in the compact factor of $\mathfrak{g}$, we prove the following:
@@
Let $x,y\in\operatorname{\mathfrak{g}}$. Consider the statements.
@@
    $\operatorname{dist}(e^{sx},e^{sy})=|sy-sx|$ for sufficiently small $s$.
@@
Then $(1)\Leftrightarrow(2)\Rightarrow(3)$. Moreover
@@
    If the norm is smooth then all the conditions are equivalent.
@@
    If the norm is strictly convex then all the conditions are equivalent to $[x,y]=0$.
We show an example that shows that the theorem is sharp. The geometry of homogeneous spaces of these groups with bi-invariant distance is part of our forthcoming research.
@@ 2.1 Lie groups with $\operatorname{Ad}$-invariant Finsler norms
The first inclusion is clear, the second one comes from Cartan’s criterion of solvability [17, page 20]: an ideal $\mathfrak{a}$ is solvable if and only if $\operatorname{\mathfrak{B}}(v,z)=0$ for all $v\in[\mathfrak{a},\mathfrak{a}]$ and all $z\in\mathfrak{a}$.
@@ 2.1.1 Hermitian operators, numerical range and spectrum
Being and ideal, we have that $x,y]\in\operatorname{\mathrm{Rad}(\operatorname{\mathfrak{g}})}$ also. But $\operatorname{\mathrm{Rad}(\operatorname{\mathfrak{g}})},\operatorname{\mathrm{Rad}(\operatorname{\mathfrak{g}})}]\subset\operatorname{\mathfrak{g}}^{\perp} [$9, Chapter I, $\S$5.5] hence $x,[x,y]]\in\operatorname{\mathfrak{g}}^{\perp}=Z(\operatorname{\mathfrak{g}})$. Thus $y\in Z(\operatorname{\mathfrak{g}})$ and $[x,y]=0$ implying $y=0$, a contradiction. ∎
@@
By the previous theorem, the results in [9, Chapter I, $\S$6.4] and in $\S$6.8 loc. cit., we have that $\operatorname{\mathfrak{g}}$ is a reductive Lie algebra, and we obtain the assertions for the Lie algebra from the Levi-Malcev theorem. Now since $Z(\operatorname{\mathfrak{g}})\oplus\operatorname{\mathfrak{k}}\ni z+s\mapsto e^{s}e^{z}$ is a local diffeomorphism around $1$ and $G$ is connected, we have the direct product decomposition of $G=H\times K$. Since the Lie algebra of $K$ is compact, the group $K$ is compact with finite center by H. Weyl’s theorem ([10] Chapter IX, $\S$1.4). ∎
@@
    Milnor’s results [30] apply, in particular his Lemma 7.5: $G$ is isomorphic to the product of a compact group and an additive vector group. Moreover, when given $G$ the left-invariant Riemannian metric induced by any $\operatorname{Ad}_{G}$ invariant inner product, the geodesics are one-parameter groups (this follows from Milnor’s formula for the covariant derivative $\nabla_{x}=\frac{1}{2}\operatorname{ad}x$, see page [30, page 323]) and the sectional curvature is always non-negative (Corollary 1.4, loc. cit.).
@@ 2.2 Distances and geodesics in the metric space setting
The assertions about distance and minimizing paths follow from the previous theorem and [24, Corollary 4.12]. The assertions about the inequality are proved in Theorem 4.17, loc. cit. ∎
@@
where $F$ is the holomorphic map $F(\lambda)=\frac{1-e^{-\lambda}}{\lambda}$. Recall that the spectrum of $\operatorname{ad}v$ is inside the interval $i(-\|\operatorname{ad}v\|,\|\operatorname{ad}v\|)$ (Definition 2.11 and Theorem 2.12). Then it must be $\|\operatorname{ad}v\|<\pi$ for all $v\in B$ otherwise we could find $v_{0}=t_{0}v$ such that $0\in\sigma(F(\operatorname{ad}v))=F(\sigma(\operatorname{ad}v)$, contradicting the invertibility of $D\exp_{v}$. Shrink a bit the ball considering $v\in rB$, with $0<r<1$, then $\|\operatorname{ad}v\|\leq M(r)<\pi$ for all $v\in rB$. Since $|\operatorname{ad}v(x)|\leq\|\operatorname{ad}v\||w|$, and with the same proof as in [2, Proposition 4.6] we obtain
@@
$\operatorname{dist}(e^{rx},e^{ry})=|r(y-x)+\frac{1}{48}r^{3}[x+y,[x,y]]+\frac{1}{16}r^{3}[y-x,[x,y]]+O(r^{4})|.$
@@ 3 Curvature
Our notions and definitions are motivated by the following remark by J. Milnor in his Morse Theore lecture notes [29, p.101]: consider an observer at $p$ looking in the direction of the unit vector $U$ towards a point $q=\exp_{p}(rU)$. A small line segment at $q$ with length $L$, pointed in a direction corresponding to the unit vector $V\in T_{p}M$, would appear to the observer as the line segment of length
@@ 3.1 Faces of the sphere, smoothness and convexity
    The norm $|\cdot|$ is Gâteaux differentiable at $x\neq 0$ if and only if it is Fréchet differentiable (this follows from Šmulian Lemma, see [16, Lemma 8.4] for instance). Therefore we simply say that the norm is smooth when this happens for any $x\neq 0$; in that case the norm function is in fact $C^{1}$ away from $x=0$ [16, Corollary 8.5].
@@
    The norm is smooth if and only if the dual norm is strictly convex if and only if there is a unique functional norming each $x\neq 0$ [16, Lemma 8.4 and Fact 8.12]. Another proof of this last assertion can be derived from

### v2

@@ Abstract
We prove that a Lie group $G$ admitting a bi-invariant distance is necessarily of the form $G=H\times K$, where $H$ is a connected abelian Lie group and $K$ is a connected compact Lie group with discrete centre. Moreover, the distance arises from a unique $\operatorname{Ad}$-invariant Finsler norm on the Lie algebra $\mathfrak{g}=Lie(G)$ as the infimum of lengths of rectifiable paths. The distance-minimising paths are left or right translates of one-parameter subgroups (though not uniquely so when the norm fails to be smooth or strictly convex). Following Milnor, we define a notion of sectional curvature $\sec(\pi)$ for any $2$-plane $\pi\subset\operatorname{\mathfrak{g}}$; this curvature is non-negative and vanishes precisely on abelian planes. When the norm is strictly convex, $\sec(\pi)=0$ if and only if $\pi$ is abelian. We obtain finer characterisations of vanishing curvature in the non-strictly-convex case. [^4] Keywords and phrases: bi-invariant metric, curvature, Finsler metric, geodesic, Lie group, one-parameter group, skew-Hermitian matrix, unitary group, unitarily invariant norm
@@ 1 Introduction
The study of Lie groups equipped with invariant metrics has a rich history, dating back to foundational works in Riemannian geometry and continuing through extensions to more general metric structures. In the Riemannian setting, Milnor’s seminal 1976 paper [25] characterized connected Lie groups admitting bi-invariant Riemannian metrics as direct products of compact semisimple Lie groups and abelian vector groups, with non-negative sectional curvature and geodesics realized as one-parameter subgroups. This result built on earlier observations that compact Lie groups always admit bi-invariant metrics (via averaging), while non-compact ones do so only under restrictive conditions. Milnor’s curvature formulas, involving the Lie bracket, highlighted the interplay between algebraic structure and geometric properties, such as the covariant derivative being half the adjoint representation.
Subsequent research extended these ideas beyond the Riemannian case. For instance, semi-Riemannian (including Lorentzian) metrics on Lie groups were explored, generalizing Milnor’s curvature computations to indefinite signatures, with applications to relativity and pseudo-Riemannian symmetric spaces (see for instance the book by O’Neill [28]). In the context of non-negative curvature, Tapp [32] classified left-invariant metrics on low-dimensional compact Lie groups like SO(3) and U(2), showing that many arise via Cheeger’s deformation technique—shrinking along subgroup chains—while others, such as certain ”twisted” metrics on U(2), do not. These classifications reveal a broader landscape of metrics with non-negative sectional curvature, often strictly containing the bi-invariant ones.
More recently, attention has turned to (smooth) Finsler metrics, which generalize Riemannian structures by allowing non-quadratic norms on tangent spaces. Latifi and Toomanian [16] studied bi-invariant Finsler metrics on Lie groups, proving that absolutely homogeneous ones render the group a symmetric Finsler space of Berwald type, with geodesics coinciding with those of an underlying bi-invariant Riemannian metric. They derived an explicit flag curvature formula, extending Milnor’s sectional curvature to the Finsler setting, and showed that such metrics exist on non-simple or higher-rank compact Lie groups but reduce to Riemannian ones on simple rank-one groups. This aligns with Berestovskii’s results [5] on homogeneous manifolds with intrinsic metrics, which imply that bi-invariant distances on locally contractible topological groups arise from Ad-invariant Finsler norms on their Lie algebras.
In our prior work [21] we established properties of Ad-invariant Finsler norms on skew-Hermitian matrices, including geometry of spheres and subdifferentials. Hofer’s metric, a Finsler-type distance arising from symplectic geometry, has been studied on compact Lie groups [20], revealing minimal geodesics and distance bounds that turned out to be helpful insights in this systematization project.
This paper builds on these foundations, studying the geometry of Lie groups with bi-invariant distances in the general Finsler context, without assuming smoothness or full homogeneity of the norm. We consider an Ad-invariant Finsler norm on the Lie algebra $\operatorname{\mathfrak{g}}=Lie(G)$, a subadditive, positively homogeneous map $|\cdot|:\operatorname{\mathfrak{g}}\to[0,\infty)$ that is non-degenerate ($|x|=0\Leftrightarrow x=0$). Such a norm induces the vector-space topology on $\operatorname{\mathfrak{g}}$ and, via left/right translation, a bi-invariant (possibly non-reversible) distance on $G$ given by the infimum of lengths of rectifiable paths. Conversely, any bi-invariant distance on a connected locally compact locally contractible topological group arises in this way from a unique Ad-invariant Finsler norm (Theorem A below).
Our aims are twofold: to characterize Lie groups admitting bi-invariant distances and to develop a notion of curvature for such geometries, particularly characterizing flat 2-planes. We adopt the metric geometry viewpoint, where curvature measures distance distortion relative to model spaces. Adapting Milnor’s ideas [24, p. 101], we define curvature along pairs of directions in $\operatorname{\mathfrak{g}}$ (see also [2] and [17] for a hyperbolic geometry analogue). In terms of complexifications, if $G$ is a compact complex Lie group with Lie algebra $\operatorname{\mathfrak{g}}$ and $\mathfrak{k}$ its real compact form ($\operatorname{\mathfrak{g}}=\mathfrak{k}\oplus i\mathfrak{k}$), our work examines the geometry of $G_{\mathbb{R}}\simeq\mathfrak{k}$ with a (non-Riemannian) bi-invariant distance, complementing studies of homogeneous spaces like $G/G_{\mathbb{R}}\simeq i\mathfrak{k}$.
What follows is an overview of the organization and key results. Motivated by Milnor’s Riemannian characterization, we address whether a similar decomposition holds for general bi-invariant distances. Affirmatively, the main result of Section 2 is:
@@
If $(G,d)$ is a connected, locally compact, locally contractible topological group with a bi-invariant intrinsic distance, then $G=H\times K$ is the product of a connected abelian Lie group $H$ and a connected compact Lie group $K$ with discrete center. In particular, $Lie(G)=Z(\operatorname{\mathfrak{g}})\oplus\mathfrak{k}$ where $\mathfrak{k}$ is a compact Lie algebra.
Locally, the distance between $e^{sx}$ and $e^{sy}$ ($x,y\in\operatorname{\mathfrak{g}}$, $s\in\mathbb{R}$) equals the norm of the Baker–Campbell–Hausdorff series of $e^{sy}e^{-sx}$,
$d(e^{sx},e^{sy})=|BCH(sy,-sx)|=|sy-sx-\frac{s}{2}[y,x]+\dots|.$
This was established in prior work [19], refined here using distance invariance (Equation (8)). If $[x,y]=0$, then $d(e^{sx},e^{sy})=|sy-sx|$ for small $s$. To explore this relation deeply, Section 3.3 adapts Milnor’s approach [24, p. 101], defining sectional curvature for a 2-plane $\pi\subset\operatorname{\mathfrak{g}}$ as
$S(x,y)=6|y-x|^{2}\lim_{r\to 0^{+}}\frac{r|y-x|-d(e^{rx},e^{ry})}{r^{2}d(e^{rx},e^{ry})}.$
Using BCH expansions, we express $S(x,y)$ in metric terms and prove:
@@
For $x,y\in\operatorname{\mathfrak{g}}$ and any Ad-invariant norm on $\operatorname{\mathfrak{g}}$,
$S(x,y)=6|y-x|\lim_{r\to 0^{+}}\frac{d(e^{r^{2}x},e^{r^{2}y})-rd(e^{rx},e^{ry})}{r^{4}}=-\frac{|y-x|}{4}\max_{\varphi\in N_{y-x}}\varphi([x,[x,y-x]])\geq 0,$
where $N_{y-x}$ is the set of unit-norm functionals with $\varphi(y-x)=|y-x|$.
Milnor observed that for Riemannian metrics with normalized $x,y$, $S(x,x+y)=\langle R(x,y)y,x\rangle$, where $R$ is the curvature tensor. Theorem B characterizes this via norming functionals (subdifferential of the norm), generalizing to Finsler cases.
In the final section, we analyze flatness using root systems in $\operatorname{\mathfrak{g}}$’s compact factor, proving in Theorem 3.28:
@@
Let $x,y\in\operatorname{\mathfrak{g}}$. Consider:
@@
    $d(e^{sx},e^{sy})=|sy-sx|$ for sufficiently small $s$.
@@
Then (1) $\Leftrightarrow$ (2) $\Rightarrow$ (3). Moreover:
@@
    If the norm is smooth, all conditions are equivalent.
@@
    If the norm is strictly convex, all are equivalent to $[x,y]=0$.
We provide an example showing sharpness, drawn from orthogonal/unitary groups with spectral or trace norms (wich are neither smooth nor strictly convex). The distance condition is related to a certain path contained in a face of the sphere of the Finsler norm, as explained in Section 3.3. We finish the paper with a convenient definition of sectional curvature $sec(\pi)$, $\pi\subset\operatorname{\mathfrak{g}}$, that takes into account the non-symmetry of the spheres (and extends the Riemannian defintion). We show that under reasonable normalization hypothesis, in this setting one always has $0\leq sec(\pi)\leq 1$.
The geometry of the homogeneous spaces of groups with bi-invariant distances is part of our forthcoming research, based on these results.
@@ 2.1 Lie groups with $\operatorname{Ad}$-invariant Finsler norms
The first inclusion is clear, the second one comes from Cartan’s criterion of solvability [14, page 20]: an ideal $\mathfrak{a}$ is solvable if and only if $\operatorname{\mathfrak{B}}(v,z)=0$ for all $v\in[\mathfrak{a},\mathfrak{a}]$ and all $z\in\mathfrak{a}$.
@@ 2.1.1 Hermitian operators, numerical range and spectrum
Being and ideal, we have that $x,y]\in\operatorname{\mathrm{Rad}(\operatorname{\mathfrak{g}})}$ also. But $\operatorname{\mathrm{Rad}(\operatorname{\mathfrak{g}})},\operatorname{\mathrm{Rad}(\operatorname{\mathfrak{g}})}]\subset\operatorname{\mathfrak{g}}^{\perp} [$7, Chapter I, $\S$5.5] hence $x,[x,y]]\in\operatorname{\mathfrak{g}}^{\perp}=Z(\operatorname{\mathfrak{g}})$. Thus $y\in Z(\operatorname{\mathfrak{g}})$ and $[x,y]=0$ implying $y=0$, a contradiction. ∎
@@
By the previous theorem, the results in [7, Chapter I, $\S$6.4] and in $\S$6.8 loc. cit., we have that $\operatorname{\mathfrak{g}}$ is a reductive Lie algebra, and we obtain the assertions for the Lie algebra from the Levi-Malcev theorem. Now since $Z(\operatorname{\mathfrak{g}})\oplus\operatorname{\mathfrak{k}}\ni z+s\mapsto e^{s}e^{z}$ is a local diffeomorphism around $1$ and $G$ is connected, we have the direct product decomposition of $G=H\times K$. Since the Lie algebra of $K$ is compact, the group $K$ is compact with finite center by H. Weyl’s theorem ([8] Chapter IX, $\S$1.4). ∎
@@
    Milnor’s results [25] apply, in particular his Lemma 7.5: $G$ is isomorphic to the product of a compact group and an additive vector group. Moreover, when given $G$ the left-invariant Riemannian metric induced by any $\operatorname{Ad}_{G}$ invariant inner product, the geodesics are one-parameter groups (this follows from Milnor’s formula for the covariant derivative $\nabla_{x}=\frac{1}{2}\operatorname{ad}x$, see page [25, page 323]) and the sectional curvature is always non-negative (Corollary 1.4, loc. cit.).
@@ 2.2 Distances and geodesics in the metric space setting
The assertions about distance and minimizing paths follow from the previous theorem and [19, Corollary 4.12]. The assertions about the inequality are proved in Theorem 4.17, loc. cit. ∎
@@
where $F$ is the holomorphic map $F(\lambda)=\frac{1-e^{-\lambda}}{\lambda}$. Recall that the spectrum of $\operatorname{ad}v$ is inside the interval $i(-\|\operatorname{ad}v\|,\|\operatorname{ad}v\|)$ (Definition 2.11 and Theorem 2.12). Then it must be $\|\operatorname{ad}v\|<\pi$ for all $v\in B$ otherwise we could find $v_{0}=t_{0}v$ such that $0\in\sigma(F(\operatorname{ad}v))=F(\sigma(\operatorname{ad}v)$, contradicting the invertibility of $D\exp_{v}$. Shrink a bit the ball considering $v\in rB$, with $0<r<1$, then $\|\operatorname{ad}v\|\leq M(r)<\pi$ for all $v\in rB$. Since $|\operatorname{ad}v(x)|\leq\|\operatorname{ad}v\||w|$, and with the same proof as in [1, Proposition 4.6] we obtain
@@
$\operatorname{dist}(e^{rx},e^{ry})=|r(y-x)+\frac{1}{48}r^{3}[x+y,[x,y]]+\frac{1}{16}r^{3}[y-x,[x,y]]+O(r^{4})|.$ (8)
@@ 3 Curvature
Our notions and definitions are motivated by the following remark by J. Milnor in his Morse Theore lecture notes [24, p.101]: consider an observer at $p$ looking in the direction of the unit vector $U$ towards a point $q=\exp_{p}(rU)$. A small line segment at $q$ with length $L$, pointed in a direction corresponding to the unit vector $V\in T_{p}M$, would appear to the observer as the line segment of length
@@ 3.1 Faces of the sphere, smoothness and convexity
    The norm $|\cdot|$ is Gâteaux differentiable at $x\neq 0$ if and only if it is Fréchet differentiable (this follows from Šmulian Lemma, see [13, Lemma 8.4] for instance). Therefore we simply say that the norm is smooth when this happens for any $x\neq 0$; in that case the norm function is in fact $C^{1}$ away from $x=0$ [13, Corollary 8.5].
@@
    The norm is smooth if and only if the dual norm is strictly convex if and only if there is a unique functional norming each $x\neq 0$ [13, Lemma 8.4 and Fact 8.12]. Another proof of this last assertion can be derived from